A regular drinking cup has a circular lower base 8 cm in diameter, a top lip 12 cm in diameter, and a height of 10 cm. What is its volume?
Answer by checkley77(12844)

height=10cm.
top diameter=12 cm.
bottom diameter=8 cm.
first we find the average diameter
(12+8)/2=10 cm is the average diameter.
now we find the area of a 10 in diameter circle or a radius of 5 cm.
area=pir^2
area=3.14*5^2
area=3.14*25
area=78.5 cm^2
now we multiply this area by the height of 10.
78.5*10=785 cm^3 is the volume of this cup.